OpenStudy (anonymous):

A plane traveling at 180 mph passes 400 feet directly over an observer. The plane is traveling along a path with an angle of elevation of 14 degrees. Find the distance of the plane from the observer 10 seconds after the plane has passed directly overhead.

OpenStudy (dumbcow):

 Find distance using Law of Cosines \[x^2 = 400^2 + 264^2 - 2(400)(264) \cos 104\]

OpenStudy (dumbcow):

the 264 is feet plane traveled in 10 sec \[180*\frac{5280}{3600}*10 = 264\]